# Blind Baselines Beat Membership Inference Attacks for Foundation Models

### v1

@@
Abstract
Membership inference (MI) attacks try to determine if a data sample was used to train a machine learning model. For foundation models trained on unknown Web data, MI attacks can be used to detect copyrighted training materials, measure test set contamination, or audit machine unlearning. Unfortunately, we find that evaluations of MI attacks for foundation models are flawed, because they sample members and non-members from different distributions. For 8 published MI evaluation datasets, we show that blind attacks—that distinguish the member and non-member distributions without looking at any trained model—outperform state-of-the-art MI attacks. Existing evaluations thus tell us nothing about membership leakage of a foundation model’s training data.
@@ 1 Introduction
Many foundation models Bommasani et al. (2021) such as GPT-4 OpenAI (2023), Gemini Gemini Team (2023) and DALL-E Ramesh et al. (2021) are trained on unknown data. There is great interest in methods that can determine if a piece of data was used to train these models. Such methods—called membership inference attacks Shokri et al. (2017)—have been studied for foundation models in many recent works Shi et al. (2023); Duan et al. (2024); Ko et al. (2023); Dubiński et al. (2024); Zhang et al. (2024); Meeus et al. (2023). Applications include privacy attacks Carlini et al. (2021), demonstrating the use of copyrighted material Meeus et al. (2024), detecting test data contamination Oren et al. (2023), or auditing the efficacy of methods to “unlearn” training data Bourtoule et al. (2021).
@@
Unfortunately, we show these strategies are severely flawed, and create easily distinguishable member and non-member distributions. As a special case of this flaw, concurrent work of Duan et al. (2024) and Maini et al. (2024) finds a temporal shift between members and non-members in the Wiki-MIA dataset Shi et al. (2023). We show this issue is persistent by identifying significant distribution shifts (beyond temporal shifts) in 8 MI evaluation datasets for foundation models, for both text and vision. Worse, we show that existing MI attacks perform “worse than chance” on these datasets. Specifically, we design “blind” attacks, which completely ignore the target model, and outperform all reported results from state-of-the-art MI attacks (see Table 1).
Our methods are naive: for some datasets with a temporal shift, we just apply a threshold to specific dates extracted from each sample; for other text or text-vision datasets, we build simple bag-of-words or n-gram classifiers on captions.
Current MI attacks for foundation models thus cannot be relied on, as we cannot rule out that they are (poorly) inferring membership based on data features, without extracting any actual membership leakage from the model. Future MI attacks should be evaluated on models with a clear train-test split, e.g., using the Pile Gao et al. (2020) or a random subset of DataComp Gadre et al. (2024) or DataComp-LM Li et al. (2024).
Table 1: MI evaluation datasets for foundation models are flawed. Our blind attacks distinguish members from non-members better than existing attacks, without looking at any model.
MI dataset Metric Best Reported Ours
WikiMIA1 TPR@5%FPR 43.2%2 94.4%
BookMIA1 AUC ROC 88.0%1 90.5%
Temporal Wiki3 AUC ROC 79.6%3 79.9%
Temporal arXiv3 AUC ROC 72.3%3 73.1%
ArXiv-1 month6 TPR@1%FPR 5.9%6 13.4%
Multi-Webdata4 TPR@1%FPR 40.3%4 83.5%
LAION-MI5 TPR@1%FPR 2.5%5 9.9%
Gutenberg6 TPR@1%FPR 18.8%6 59.6%
1Shi et al. (2023)   2Zhang et al. (2024)   3Duan et al. (2024)
4Ko et al. (2023)  5Dubiński et al. (2024)    6Meeus et al. (2023)
@@ 2 Background and Related Work
Membership inference attacks have been applied to LLMs Shi et al. (2023); Duan et al. (2024); Zhang et al. (2024); Meeus et al. (2023), diffusion models Dubiński et al. (2024), CLIP Ko et al. (2023), and other foundation models. The motivations for these attacks include using them as a component of a privacy attack Carlini et al. (2021), for evaluating unlearning methods Shi et al. (2023), or for data provenance tests to detect the use of copyrighted data Meeus et al. (2024). Due to the lack of a dedicated test set (and possibly even an unknown training set) for the targeted models, many of these works design custom evaluation datasets for membership inference attacks, by collecting sources of data that were likely used (respectively not used) for training.
@@
Membership inference evaluations are usually set up so that a baseline attack (which does not query the target model) achieves 50% AUC (or equal TPR and FPR). Some authors have also considered cases where the attacker has a non-uniform prior Jayaraman et al. (2020). In either case, the goal of a MI attack is to extract information from the model to beat a baseline inference without access to the model.
Many MI evaluation datasets for foundation models introduce distribution shifts between members and non-members, which allows for baseline attacks with non-trivial success. In concurrent work, Duan et al. (2024) and Maini et al. (2024) identify a temporal shift between members and non-members in one dataset—WikiMIA Shi et al. (2023)—and demonstrate that some attacks fail when temporal shifts are removed. Our work shows that this issue is much broader: virtually all evaluation sets proposed for membership inference on foundation models are flawed. We further show that existing attacks do not just exploit distribution shifts, but they do so sub-optimally and are easily beat by blind baselines. These attacks thus perform “worse than chance”.
@@ 3.1 Distribution Shifts in MI Evaluation Datasets
MI evaluations measure the ability to distinguish a model’s training set members from non-members. A baseline attack (with no knowledge of the model) should do no better than random guessing—i.e., an AUC of 50% or equal TPR and FPR. For current MI evaluations sets for foundation models, this is not the case, due to intrinsic differences between members and non-members. We discuss the most common reasons for this discrepancy below.
@@
MI evaluation sets based on a hard data cutoff for some evolving Web source (e.g., Wikipedia or arXiv) introduce a temporal shift between members and non-members. We can thus blindly distinguish members from non-members if we can answer a question of the form: “is this data from before or after 2023?”
@@ 3.2 Blind Attack Techniques
In this section, we introduce simple “blind” attack techniques to distinguish members from non-members in MI evaluations. The datasets we consider consist of either text, or images and text. For simplicity, we focus only on the text modality in both cases. We do not aim for our blind attacks to be optimal. We prioritize simple and interpretable methods to show that existing evaluation datasets suffer from large distribution shifts. Our blind attacks typically aim to achieve a high TPR at a low FPR (i.e., very confident predictions of membership for as many samples as possible).
@@ 4 Case Studies
Table 2: Extended version of Table 1, with the results from our blind attacks compared to the best reported MI attack on each evaluation set. We report the same metrics as used in prior work, and add TPR at 1% FPR if no prior results are reported.
MI dataset Metric Best attack Ours
@@
WikiMIA TPR@5%FPR 43.2% 94.4%
AUC ROC 83.9% 98.7%
BookMIA TPR@5%FPR - 60.8%
AUC ROC 88.0% 90.5%
Temporal Wiki TPR@1%FPR - 33.5%
AUC ROC 79.6% 79.9%
Temporal arXiv TPR@1%FPR - 10.8%
AUC ROC 74.5% 75.6%
ArXiv-1 month TPR@1%FPR 5.9% 13.4%
AUC ROC 67.8% 68.4%
@@
Multi-Webdata TPR@1%FPR 40.3% 83.5%
AUC ROC 81.7% 96.9%
LAION-MI TPR@1%FPR 2.5% 9.9%
Gutenberg TPR@1%FPR 18.8% 59.6%
AUC ROC 85.6% 96.4%
* Date detection    $\dagger$ Bag-of-words    $\ddagger$ Greedy rare words
We now study 8 membership inference datasets proposed for large language models and diffusion models. Table 2 summarizes our results: for each dataset, we create blind attacks that outperform existing MI attacks that have access to a trained model. We average the results from multiple runs and report for the same metric used in prior work, either AUC ROC or TPR at low FPR. Whenever possible, we use the exact datasets released by the authors to ensure that no biases are introduced. For datasets that are not publicly available (arXiv-1 month and Gutenberg Meeus et al. (2023)), we follow the specific collection steps outlined in the respective work to create similar datasets.
@@ 4.1.1 WikiMIA
The WikiMIA dataset Shi et al. (2023) selects members from Wikipedia event pages from before 01/01/2017 and non-members from after 01/01/2023. It thus serves as an MI evaluation set for any LLM trained in between those dates. The best reported prior MI attack on this evaluation set is from the Min-K%++ method of Zhang et al. (2024).
@@
We first apply our naive date detection attack. We extract all dates in the snippet and check if the latest one is from before 2023. Using this, we obtain a $52.3$% TPR at $5$% FPR. This already beats the state-of-the-art MI attack evaluated on this dataset (see Table 2). To obtain an even stronger blind attack, we train a bag-of-words classifier on 80% of the dataset, and evaluate on the remaining 20%. This classifier achieves a near perfect TPR of $94.4$%.
@@ 4.1.2 BookMIA
We train a a bag-of-words classifier on 80% of the dataset, and evaluate on the remaining 20%. As shown in Table 2, a bag-of-words classifier achieves a $90.5$% AUC ROC and beats state-of-the-art MI attacks against GPT-3.111To control for book- and author-specific features (such as character names), we repeat the experiment with no author appearing in both our classifier’s train and test sets. This still results in a significantly above chance AUC ROC of $80.3$%.
@@ 4.1.3 Temporal Wiki & arXiv
Duan et al. (2024) hypothesize that some MI attacks evaluated on Wiki-MIA may inadvertently exploit temporal shifts between members and non-members. To showcase this, they create datasets that sample members from the Wikipedia and arXiv snippets in the Pile Gao et al. (2020) and non-members from the same sources at a later date.
@@
We show that even if existing attacks do rely on some temporal features in the Temporal Wikipedia and Temporal arXiv datasets, they do so sub-optimally. As for the prior datasets, we train a bag-of-words classifier which outperforms prior MI results on these datasets. Specifically, we focus on the “2020-08” split (where non-members are taken from arXiv articles in August 2020, right after the Pile cutoff). There, our blind attack slightly outperforms the AUC and TPR at 1% FPR of the best reported MI attack on both datasets (see Table 2).
4.1.4 One-month arXiv
@@ 4.2 Biases in Data Replication and Distinguishable Tails
4.2.1 Multi-Webdata
The dataset and evaluation.
To evaluate MI attacks on diffusion models, Ko et al. (2023) collect multiple web-scale datasets of captioned images: CC3M Sharma et al. (2018), CC12M Changpinyo et al. (2021), (LAION Schuhmann et al. (2022) and MS-COCO Lin et al. (2014)). They argue that since these datasets are sampled from a common global source (i.e., captioned images from the Web), they can serve as a natural evaluation set for MI attacks. Specifically, the authors train a diffusion model on CC12M (the members), and sample the non-members from the union of the other three datasets. [^2] Wu et al. (2023) use a similar approach, and choose members and non-members from different text-vision datasets. Ko et al. (2023) evaluate the performance of three newly proposed MI attacks on target models from the ViT Dosovitskiy et al. (2020) and ResNet He et al. (2016) families trained on CC12M.
Our attack.
For simplicity, we focus only on the image captions and ignore the visual content. We train a bag-of-words binary classifier on the captions, achieving a TPR of $75.7\%$ at a $1$% FPR, far superior to the best MI attack methods evaluated on these datasets (see Table 2). It is likely that incorporating visual features into the classifier could lead to an even stronger blind attack.
@@ 4.2.2 LAION-MI
To align the two distributions, the authors first apply the same selection process that was used to select images in LAION: (1) they discard all images from multilingual LAION with an “aesthetics score” below some threshold; (2) they translate the non-member captions to English, using Facebook’s M2M100 1.2B They then train a distinguisher on top of image captions, and select a subset of the data where members and non-members are hard to distinguish on average. The authors use a PCA visualization to confirm that the two distributions closely match.
@@
Table 3: Characters with the highest distinguishing power among members and non-members for LAION-MI.
Character Members Non-members
| 0.60% 0.00%
@@
0.21% 0.00%
о 0.23% 0.03%
т 0.15% 0.03%
To boost the power of this approach, we apply our greedy selection algorithms to all n-grams of up to 5 characters. We use 80% of the data to pick the n-grams with the highest TPR-to-FPR ratio until we hit a $1\%$ FPR. At that point, we obtain a $9.9$% TPR on the evaluation set—surpassing the best MI attack.

### v2

@@
Blind Baselines Beat Membership Inference Attacks for Foundation Models
Debeshee Das, Jie Zhang, Florian Tramèr
ETH Zurich
Аннотация
Membership inference (MI) attacks try to determine if a data sample was used to train a machine learning model. For foundation models trained on unknown Web data, MI attacks are often used to detect copyrighted training materials, measure test set contamination, or audit machine unlearning. Unfortunately, we find that evaluations of MI attacks for foundation models are flawed, because they sample members and non-members from different distributions. For 8 published MI evaluation datasets, we show that blind attacks—that distinguish the member and non-member distributions without looking at any trained model—outperform state-of-the-art MI attacks. Existing evaluations thus tell us nothing about membership leakage of a foundation model’s training data.
@@ 1 Introduction
Many foundation models Bommasani et al. (2021) such as GPT-4 OpenAI (2023), Gemini Gemini Team (2023) and DALL-E Ramesh et al. (2021) are trained on unknown data. There is great interest in methods that can determine if a piece of data was used to train these models. Such methods—called membership inference attacks Shokri et al. (2017)—have been studied for foundation models in many recent works Shi et al. (2023); Duan et al. (2024); Ko et al. (2023); Dubiński et al. (2024); Zhang et al. (2024a); Meeus et al. (2023). Applications include privacy attacks Carlini et al. (2021), demonstrating the use of copyrighted material Meeus et al. (2024b), detecting test data contamination Oren et al. (2023), or auditing the efficacy of methods to ‘‘unlearn’’ training data Bourtoule et al. (2021).
@@
Unfortunately, we show these strategies are severely flawed, and create easily distinguishable member and non-member distributions. As a special case of this flaw, concurrent works of Duan et al. (2024), Maini et al. (2024) and Meeus et al. (2024a) find a temporal shift between members and non-members in the Wiki-MIA dataset Shi et al. (2023). We show this issue is persistent by identifying significant distribution shifts (beyond temporal shifts) in 8 MI evaluation datasets for foundation models, for both text and vision. Worse, we show that existing MI attacks perform ‘‘worse than chance’’ on these datasets. Specifically, we design ‘‘blind’’ attacks, which completely ignore the target model, and outperform all reported results from state-of-the-art MI attacks (see Table 1).
Our methods are naive: for some datasets with a temporal shift, we apply a threshold to specific dates extracted from each sample; for other text or text-vision datasets, we build simple bag-of-words or n-gram classifiers on captions. We find that such simple methods work even for MI evaluation datasets that were explicitly designed to remove distribution shifts between members and non-members. Our work shows that removing biases in a post-hoc fashion is highly brittle.
Current MI attacks for foundation models thus cannot be relied on, as we cannot rule out that they are (poorly) inferring membership based on data features, without extracting any actual membership leakage from the model. In addition, studies that rely on the MI evaluations on these datasets, also cannot be trusted. For example, the authors of Panaitescu-Liess et al. (2024) investigate the effect of watermarking on MI attacks and evaluate it on datasets that we show to have clear distribution shifts. Future MI attacks should be evaluated on models with a clear train-test split, e.g., using the Pile Gao et al. (2020) or a random subset of DataComp Gadre et al. (2024) or DataComp-LM Li et al. (2024).
Таблица 1: MI evaluation datasets for foundation models are flawed. Our blind attacks distinguish members from non-members better than existing attacks, without looking at any model.
MI dataset Metric Best Reported (%) Ours (%)
WikiMIA TPR@5%FPR 43.2 94.7
BookMIA AUC ROC 88.0 91.4
Temporal Wiki AUC ROC 79.6 79.9
Temporal ArXiv AUC ROC 74.5 75.3
ArXiv (all vs 1 month) TPR@1%FPR 5.9 10.6
ArXiv (1 month vs 1 month) TPR@1%FPR 2.5 2.7
LAION-MI TPR@1%FPR 2.5 8.9
Gutenberg TPR@1%FPR 18.8 55.1
@@ 2 Background and Related Work
Membership inference attacks have been applied to LLMs Shi et al. (2023); Duan et al. (2024); Zhang et al. (2024a); Meeus et al. (2023), diffusion models Dubiński et al. (2024), CLIP Ko et al. (2023), and other foundation models. The motivations for these attacks include using them as a component of a privacy attack Carlini et al. (2021), for evaluating unlearning methods Shi et al. (2023), or to detect the use of copyrighted data Meeus et al. (2024b). Due to the lack of a dedicated test set (and possibly even an unknown training set) for the targeted models, many of these works design custom evaluation datasets for membership inference attacks, by collecting sources of data that were likely used (respectively not used) for training.
@@
Membership inference evaluations are usually set up so that a baseline attack (which does not query the target model) achieves 50% AUC (or equal TPR and FPR). Some authors have also considered cases where the attacker has a non-uniform prior Jayaraman et al. (2020). In either case, the goal of an MI attack is to extract information from the model to beat a baseline inference that does not have access to the model.
Many MI evaluation datasets for foundation models introduce distribution shifts between members and non-members, which allows for baseline attacks with non-trivial success. In concurrent work, Duan et al. (2024) and Maini et al. (2024) identify a temporal shift between members and non-members in one dataset—WikiMIA Shi et al. (2023)—and demonstrate that some attacks fail when temporal shifts are removed. Our work shows that this issue is much broader: virtually all evaluation sets proposed for membership inference on foundation models are flawed. We further show that existing attacks do not just exploit distribution shifts, but they do so sub-optimally and are easily beaten by blind baselines. These attacks thus perform ‘‘worse than chance’’.
@@ 3.1 Distribution Shifts in MI Evaluation Datasets
MI evaluations measure the ability to distinguish a model’s training set members from non-members. A baseline attack (with no knowledge of the model) should do no better than random guessing—i.e., an AUC of $50\%$ or equal TPR and FPR. For current MI evaluation datasets for foundation models, this is not the case, due to intrinsic differences between members and non-members. We discuss the most common reasons for this discrepancy below.
@@
MI evaluation sets based on a hard data cutoff for some evolving Web source (e.g., Wikipedia or arXiv) introduce a temporal shift between members and non-members. We can thus blindly distinguish members from non-members if we can answer a question of the form: ‘‘is this data from before or after 2023?’’
@@ 3.2 Blind Attack Techniques
In this section, we introduce simple ‘‘blind’’ attack techniques to distinguish members from non-members in MI evaluations. The datasets we consider consist of either text, or images and text. For simplicity, we focus only on the text modality in both cases. We do not aim for our blind attacks to be optimal. We prioritize simple and interpretable methods to show that existing evaluation datasets suffer from large distribution shifts. Our blind attacks typically aim to achieve a high TPR at a low FPR (i.e., very confident predictions of membership for as many samples as possible).
@@ 4 Case Studies
We now study 8 membership inference datasets proposed for large language models and diffusion models. Table 2 summarizes our results: for each dataset, we create blind attacks that outperform the best existing MI attacks that have access to a trained model. We average the results from multiple runs and report for the same metric used in prior work, either AUC ROC or TPR at low FPR. Whenever possible, we use the exact datasets released by the authors to ensure that no biases are introduced. For datasets that are not publicly available (arXiv-1 month and Gutenberg Meeus et al. (2023)), we follow the specific collection steps outlined in the respective work to create similar datasets.
Таблица 2: Extended version of Table 1, with the results from our blind attacks compared to the best reported MI attack on each evaluation set. We report the same metrics as used in prior work, and add TPR at 1% FPR if no prior results are reported.
Section MI Dataset Metric Best Attack  (%) Ours  (%)
@@
 4.1.1 WikiMIA TPR@5%FPR Zhang et al. (2024a) 43.2 Bag of Words 94.7
AUC ROC Zhang et al. (2024a) 83.9 Bag of Words 99.0
4.1.2 BookMIA TPR@5%FPR Zhang et al. (2024b) 33.6 Bag of Words 64.5
AUC ROC Shi et al. (2023) 88.0 Bag of Words 91.4
4.1.3 Temporal Wiki TPR@1%FPR - - Greedy 36.5
AUC ROC Duan et al. (2024) 79.6 Greedy 79.9
4.1.3 Temporal Arxiv TPR@1%FPR - - Bag of Words 9.1
AUC ROC Duan et al. (2024) 74.5 Bag of Words 75.3
4.1.4 Arxiv TPR@1%FPR Meeus et al. (2023) 5.9 Date Detection 10.6
(all vs 1 mo) AUC ROC Meeus et al. (2023) 67.8 Date Detection 72.3
4.1.5 Arxiv TPR@1%FPR Meeus et al. (2024a) 2.5 Greedy 2.7
(1 mo vs 1 mo)
@@
4.2.1 LAION-MI TPR@1%FPR Dubiński et al. (2024) 2.5 Greedy 8.9
4.2.2 Gutenberg TPR@1%FPR Meeus et al. (2023) 18.8 Greedy 55.1
AUC ROC Meeus et al. (2023) 85.6 Bag of Words 96.1
@@ 4.1.1 WikiMIA
The WikiMIA dataset Shi et al. (2023) selects members from Wikipedia event pages from before 01/01/2017 and non-members from after 01/01/2023. It thus serves as an MI evaluation set for any LLM trained in between those dates. The best reported prior MI attack on this evaluation set is from the Min-K%++ method of Zhang et al. (2024a).
@@
We first apply our naive date detection attack. We extract all dates in the snippet and check if the latest one is from before 2023. Using this, we obtain a $52.3$% TPR at $5$% FPR. This already beats the state-of-the-art MI attack evaluated on this dataset (see Table 2). To obtain an even stronger blind attack, we train a bag-of-words classifier on 80% of the dataset, and evaluate on the remaining 20%. This classifier achieves a near perfect TPR of $94.7$%. (See Appendix 1 for a visualization of the clear distribution shift in the WikiMIA dataset).
@@ 4.1.2 BookMIA
Duarte et al. (2024) extended this dataset to propose the BookTection dataset. Based on their construction, we expected this dataset to suffer from the same drawbacks as BookMIA and thus do not include it in our study.
@@
We train a a bag-of-words classifier on 80% of the dataset, and evaluate on the remaining 20%. As shown in Table 2, a bag-of-words classifier achieves a $91.4$% AUC ROC and beats state-of-the-art MI attacks against GPT-3.111To control for book- and author-specific features (such as character names), we repeat the experiment with no author appearing in both our classifier’s train and test sets. This still results in a significantly above chance AUC ROC of $80.3$%.
@@ 4.1.3 Temporal Wiki & arXiv
Duan et al. (2024) hypothesize that some MI attacks evaluated on Wiki-MIA may inadvertently exploit temporal shifts between members and non-members. To showcase this, they create datasets that sample members from the Wikipedia and arXiv snippets in the Pile Gao et al. (2020) and non-members from the same sources at a later date. [^2] Duarte et al. (2024) propose a similar ArxivTection dataset constructed using the same temporal principle.
@@
We show that even if existing attacks do rely on some temporal features in the Temporal Wikipedia and Temporal arXiv datasets, they do so sub-optimally. We train a bag-of-words classifier which outperforms prior MI results on these datasets. Specifically, we focus on the ‘‘2020-08’’ split (where non-members are taken from arXiv articles in August 2020, right after the Pile cutoff). There, our blind attack slightly outperforms the AUC and TPR at 1% FPR of the best reported MI attack on both datasets (see Table 2).
4.1.4 ArXiv (all vs one month)
@@
4.1.5 ArXiv (one month vs one month)
The dataset and evaluation.
In concurrent work, Meeus et al. (2024a) also note the issue of a temporal shift in their previous dataset we discussed above. To remedy this, they build a new MI evaluation set where the non-members are still arXiv articles from March 2023, but the members are now limited to articles from February 2023 (i.e., right before the cutoff date of the RedPajama dataset Together (2023)). Unfortunately, this dataset is hence suitable for evaluating MI attacks on a very narrow subset of data. The hope is that this aggressive filtering can remove the temporal shift altogether. And indeed, existing MI attacks fare poorly on this split of members and non-members: the authors show that their best MI attack achieves a TPR of $2.5\%$ at an FPR of $1\%$.
Our attack.
Even a one month gap could result in a rather significant distribution shift if some research topics are emerging or disappearing from the mainstream. As an illustrative example, if we wanted to show non-membership, some topics provide a very strong signal: we find that articles from March 2023 are about 25×25$\times$25 × more likely to contain the terms ‘‘GPT-4’’ or ‘‘iccvfinalcopy’’ than articles from February 2023.333GPT-4 was released on March 14th 2023. The submission deadline for ICCV 2023 was on March 8th 2023.
Biases in the opposite direction are not as clear-cut, but still exist. Using $80\%$ of the data, we greedily select individual words that have the highest TPR-to-FPR ratio on this data (top examples include ‘‘pre-registered’’, ‘‘speaker-dependent’’ or ‘‘MMoE’’). We then predict that a sample from the held-out set is a member if and only if it contains one of these words. This yields a blind TPR of $2.7\%$ at an FPR of $1\%$, slightly better than the best MI attack that has access to a trained model. While this TPR is small, it does show that temporal biases are persistent even in simple word choices.
@@ 4.2.1 LAION-MI
To align the two distributions, the authors first apply the same selection process that was used to select images in LAION: (1) they discard all images from multilingual LAION with an ‘‘aesthetics score’’ below some threshold; (2) they translate the non-member captions to English, using Facebook’s M2M100 1.2B. They then train a distinguisher on the image captions, and select a subset of the data where members and non-members are hard to distinguish on average. The authors use a PCA visualization to confirm that the two distributions closely match.
@@
Таблица 3: Characters with the highest distinguishing power among members and non-members for LAION-MI.
Character Members (%) Non-members (%)
| 0.60 0.00
@@
0.21 0.00
о 0.23 0.03
т 0.15 0.03
To boost the power of this approach, we apply our greedy selection algorithms to all n-grams of up to 5 characters. We use $80\%$ of the data to pick the n-grams with the highest TPR-to-FPR ratio until we hit a $1\%$ FPR. At that point, we obtain an $8.9$% TPR on the evaluation set—surpassing the best MI attack.